# Video Diffusion Transformers Quantization

### v2

@@ 1 Introduction
Recent advances in diffusion transformers (DiTs) (Peebles & Xie, 2023) have revolutionized video generation (Singer et al., 2023), enabling high-fidelity synthesis through iterative denoising processes. Innovations such as Sora’s unified framework (OpenAI, 2024) and SkyReels-V2’s infinite-length generation (Chen et al., 2025a) demonstrate enhanced controllability, while efficiency gains emerge from quantization (Zhao et al., 2025), sparse attention mechanisms (Xi et al., 2025; Zhang et al., 2025) and cache techniques (Wu et al., 2025; Liu et al., 2024a). The emergence of large-scale models like HunyuanVideo (Kong et al., 2024) has further improved video generation quality with its causal 3D VAE architecture and full-attention mechanisms, achieving film-grade quality and physics-aware generation. These developments collectively demonstrate enhanced controllability in temporal coherence and resolution, though challenges persist in computational efficiency.
Prior work in diffusion model quantization has approached these challenges through two distinct paradigms. Quantization-Aware Training (QAT) requires full model fine-tuning but achieves superior low-bit performance. Ter-DiT (Lu et al., 2024) introduces RMSNorm-enhanced adaLN modules for stable ternary training of Diffusion Transformers. In contrast, Post-Training Quantization (PTQ) methods (Li* et al., 2025; Chen et al., 2025b; Zhao et al., 2025) offer deployment-friendly solutions without retraining. SVDQuant (Li* et al., 2025) employs low-rank SVD decomposition to absorb outliers into 16-bit branches for 4-bit weight/activation quantization on FLUX.1 models (Labs, 2024). ViDiT-Q (Zhao et al., 2025) introduces a set of techniques for DiT-based generative models and achieves negligible performance drop under W8A8 quantization. These approaches demonstrate the trade-off between QAT’s higher accuracy and PTQ’s faster deployment in DiT compression.
Although PTQ offers a plug-and-play alternative, existing PTQ methods still face two critical limitations. First, the calibration-based pre-scaling adopted by most video-specific quantization techniques (Zhao et al., 2025; Wu et al., 2024; Chen et al., 2025b) not only demands extensive calibration time but also struggles to adapt to timestep-dependent scale variations in DiTs. Second, aggressive W4A4 quantization results in significant performance degradation (Fig. 1), with VBench metrics dropping by 27.5% to 61.3%. Our analysis reveals three key insights to overcome these limitations: (i) Weights exhibit Gaussian-like distributions (Fig. 3), making fixed quantization ranges suboptimal for preserving critical parameters. (ii) Activation scales vary significantly across denoising timesteps, necessitating dynamic rather than static quantization strategies. (iii) Latent feature variations exist across diffenrent denoising timesteps, enabling the possibility of adaptive bit-width allocation during online inference.
@@ 2.1 Diffusion Transformers (DiT)
Diffusion Transformers (DiTs) (Peebles & Xie, 2023) represent a paradigm shift in generative modeling, replacing the traditional U-Net (Saharia et al., 2022; Blattmann et al., 2023; Ramesh et al., 2022) backbone with transformer architectures. This breakthrough was attributed to the self-attention mechanism’s ability to model long-range dependencies (Vaswani et al., 2023), proving especially beneficial for high-resolution synthesis.
The success of DiTs in image generation quickly extended to video synthesis. Latte (Ma et al., 2025) pioneered this transition by adapting the DiT framework for temporal modeling, achieving unprecedented coherence in text-to-video generation. This was followed by Sora (OpenAI, 2024), which scaled DiTs to massive parameter counts and dataset sizes, demonstrating remarkable capabilities in long video generation with complex dynamics. Open-source implementations like HunyuanVideo (Kong et al., 2024), Open-Sora (Zheng et al., 2024; Peng et al., 2025) and CogvideoX (Yang et al., 2024) further democratized these advancements, making them widely accessible.
Despite their impressive results, DiTs inherit fundamental challenges from both transformer architectures and diffusion processes. Attention’s quadratic complexity is particularly problematic for video generation, where sequence length scales with spatial resolution and frame count. Moreover, the iterative nature of diffusion models requires multiple forward passes through the entire architecture (typically 50-100 steps), with each step processing increasingly refined features. These limitations have driven efforts to boost DiTs’ efficiency, such as caching (Wu et al., 2025; Liu et al., 2024a), and quantization (Li* et al., 2025; Chen et al., 2025b; Zhao et al., 2025).
@@ 2.2 Model Quantization
Model quantization (Xiao et al., 2023; Li* et al., 2025; Liu et al., 2024b; Jacob et al., 2018a; Li et al., 2023) has emerged as a fundamental technique for compressing deep neural networks by converting full-precision parameters into low-bit representations, significantly reducing memory footprint and computational costs. Post-Training Quantization (PTQ) (Li et al., 2025; Frantar et al., 2023; Yan et al., 2025) has proven particularly effective as it compresses pre-trained models without requiring extensive retraining.
Significant progress has been made to address quantization challenges in transformers (Vaswani et al., 2017). SmoothQuant (Xiao et al., 2023) introduces channel-wise scaling to balance weight and activation quantization difficulty, while Quarot (Ashkboos et al., 2024) employs orthogonal matrix rotations to distribute values evenly across channels. For diffusion models, quantization presents unique challenges due to their time-dependent nature. Q-Diffusion (Li et al., 2023) and PTQ4DM (Shang et al., 2023) address this by collecting timestep-wise activation statistics to determine optimal quantization parameters. Q-DiT (Chen et al., 2025b) tackles channel-wise imbalance via customized quantization parameters. Subsequent work such as SVDQuant (Li* et al., 2025) introduces low-rank branches to handle activation outliers, enabling 4-bit quantization without performance degradation. For video generation, ViDiT-Q (Zhao et al., 2025) achieves lossless W8A8 quantization.
Despite these advances, text-to-video generation with quantized DiTs still suffers from two persistent limitations. First, offline calibration for activation pre-scaling in existing video quantization methods (Zhao et al., 2025; Wu et al., 2024; Chen et al., 2025b) imposes heavy computational burdens and struggles to adapt to large activation vatiants. Second, pushing quantization precision below 8 bits for activations (W4A4) triggers significant quality degradation. Our framework overcomes these obstacles through three key innovations presented in Sec. 3.
@@ 3 Method
Overview. As shown in Fig. 2, Bounded-init Grid Refinement (BGR) iteratively refines quantization grid upon progressive tightening bounds, which preserves Gaussian-distributed DiT weights with minimal error (Sec. 3.1). Auto-scaling Rotated Quantization (ARQ) eliminates calibration overhead by jointly optimizing rotation and online scaling for activation outliers, which also offers greater flexibility for timestep-dependent scale variants (Sec. 3.2). Additionally, -Guided Bit Switching further allocates bit-widths adaptively across timesteps by tracking feature evolution (Sec. 3.3). The synergy of these techniques achieves lower bit-widths (e.g., W4A4 and W4A6) with negligible quality degradation, overcoming limitations of prior quantization approaches.
@@ 3.1 Bounded-init Grid Refinement (BGR)
For weight matrix , vanilla MinMax (per-channel) quantization (Jacob et al., 2018b) directly adopts the extreme values for range calculation through Quant/DeQuant processes:
@@
Notably, the heuristic solution from MinMax algorithm provides an initialization for this optimization problem. As post-training quantization aims to avoid computationally expensive process (e.g., gradient descent), we seek a closed-form solution for the quantization objective. Assuming we have obtained the initial state of , and , our goal is to refine one or all of them to reduce quantization error. Refining them simultaneously is highly challenging, instead, we can refine one parameter while fixing the other two. For instance, we first fix and , then derive a better solution for by solving the least squares problem for Eq. 3:
@@
notably, this refinement of aligns with its definition in Eq. 1.
This sequential refinement procedure can be further extended to multiple rounds and combined with various initialization algorithms. Beyond MinMax, search bound methods (Liu et al., 2024b; Shao et al., 2023) also serve as valid initial strategies, which exclude outliers by progressively clipping full-precision weights:
where and denote the shrink step sizes for lower and upper bounds, respectively. Since search bound algorithm provides a better initial state for quantization step-size and zero-point, our grid refinement algorithm is built on this initialization (see Alg. 1).
Our experiments in Fig. 3 quantitatively show a substantial 86% reduction in quantization error, effectively preserving the critical parameters within high-density regions. These gains incur no additional inference-time computational overhead, making BGR more accurate than previous approaches.
@@ 3.2 Auto-scaling Rotated Quantization (ARQ)
Pre-scaling methods (Xiao et al., 2023; Zhao et al., 2025; Wu et al., 2024) introduce a channel-wise mask to alleviate quantization difficulty by transferring it from activations to weights:
where is a hyperparameter that controls the degree of difficulty mitigation. Moreover, these methods typically require a calibration dataset to compute scaling factors offline. When applied to DiT models (which involve 50 denoising steps (Xiao et al., 2023; Zhao et al., 2025)), the calibration set often fails to capture the full dynamic range of activation distributions across all timesteps, introducing substantial quantization errors and potential bias.
Rotation-based methods (Ashkboos et al., 2024; Liu et al., 2025; Zhao et al., 2025) employ an orthogonal rotation matrix satisfying and . While multiplying by can effectively suppress large outliers, the rotation operation may inadvertently amplify certain activation values. This risks introducing new quantization errors in the transformed space, as shown in Fig. 4.
To address these challenges, we propose Auto-scaling Rotated Quantization (ARQ) for activation quantization. Our approach combines the strengths of both rotation-based and scaling-based methods while overcoming their individual limitations. We first incorporate Hadamard matrix multiplication (Tseng et al., 2024) on both sides of activations and weights to preserve computational invariance: . Hadamard rotation is computed by fast Hadamard transform, which introduces marginal overhead to the inference latency. Subsequently, per-channel scaling factors are computed online and applied directly to activations (rather than being transferred to weights):
@@
We analyze how ARQ addresses the massive outliers and channel-wise outliers inherent in DiT activations. For massive outliers, rotation redistributes these extreme values across multiple channels. For channel-wise outliers, the post-rotation online scaling reinforces channel-wise consistency and mitigates the limitation of rotation shown in Fig. 4. Furthermore, our online scaling strategy eliminates reliance on offline calibration datasets. For DiT models, where activation distributions vary drastically across denoising steps, this runtime adaptation ensures optimal quantization parameters at every timestep, sustaining generation quality throughout the entire diffusion process.
@@
The denoising process in video DiTs exhibits non-uniform feature evolution across timesteps. Uniform quantization wastes computational resources on these redundant timesteps while potentially compromising quality during critical transformation phases. Existing adaptive quantization methods either depend on expensive calibration analysis (Zhao et al., 2025) or adopt static timestep segmentation (Wu et al., 2024). Neither of them can capture dynamic feature changes with respect to different input prompts in video generation.
@@
To further enhance video generation quality, we propose -Guided Bit Switching, tailored to input characteristics. Prior work (Kahatapitiya et al., 2024; Liu et al., 2024a; Wu et al., 2025; So et al., 2024; Ma et al., 2024) found that denoising process of DiTs contains redundant timesteps, where latent feature changes are marginal. For these redundant timesteps, we can apply lower bit-width quantization across the entire model. In contrast, we use higher precision for critical timesteps with significant feature transformations. Specifically, our mixed-precision mechanism operates as follows:
@@
where denotes the most recent timestep at which cumulative error tracking is reset, and represents the model’s output features at timestep . Our algorithm works by continuously monitoring normalized L1 distances between successive outputs. When cumulative feature changes remain below the threshold , we apply -bit quantization, indicating marginal feature evolution can tolerate lower precision. Once the accumulated error exceeds , we switch to -bit quantization to preserve critical details, while simultaneously resetting the cumulative error counter to zero. This adaptive approach optimizes bit allocation while incurring negligible additional inference overhead, with the threshold acting as a hyperparameter to balance performance and efficiency.
| Method |
| HunyuanVideo (Kong et al., 2024) | 16/16 | 62.53 | 64.78 | 25.86 | 42.81 | 97.01 | 96.05 | 51.39 | 99.30 | ||||||||||||||||||
| MinMax (Jacob et al., 2018b) | 4/8 | 59.44 | 60.62 | 25.78 | 36.41 | 97.61 | 95.83 | 52.78 | 98.89 | ||||||||||||||||||
| SmoothQuant (Xiao et al., 2023) | 4/8 | 60.50 | 64.47 | 25.56 | 28.85 | 97.72 | 96.29 | 51.39 | 99.05 | ||||||||||||||||||
| Quarot (Ashkboos et al., 2024) | 4/8 | 58.80 | 56.86 | 25.33 | 34.30 | 98.10 | 95.72 | 55.56 | 99.03 | ||||||||||||||||||
| ViDiT-Q (Zhao et al., 2025) | 4/8 | 57.01 | 59.74 | 24.77 | 27.11 | 97.37 | 95.16 | 48.61 | 99.06 | ||||||||||||||||||
| -Quant | 4/6 | 62.27 | 64.22 | 25.83 | 33.07 | 97.89 | 96.57 | 58.33 | 99.05 | ||||||||||||||||||
| MinMax (Jacob et al., 2018b) | 4/4 | 24.20 | 24.78 | 4.27 | 0.00 | 98.05 | 96.27 | 0.00 | 99.03 | ||||||||||||||||||
| SmoothQuant (Xiao et al., 2023) | 4/4 | 48.41 | 59.46 | 21.09 | 7.84 | 96.72 | 94.97 | 1.39 | 98.79 | ||||||||||||||||||
| Quarot (Ashkboos et al., 2024) | 4/4 | 44.85 | 54.30 | 17.33 | 0.94 | 97.69 | 92.64 | 87.5 | 92.22 | ||||||||||||||||||
| ViDiT-Q (Zhao et al., 2025) | 4/4 | 45.36 | 40.10 | 19.66 | 7.85 | 97.19 | 97.29 | 0.00 | 99.43 | ||||||||||||||||||

### v3

@@ 1 Introduction
Recent advances in diffusion transformers (DiTs)Β (Peebles and Xie, 2023) have revolutionized video generationΒ (Singer et al., 2023), enabling high-fidelity synthesis through iterative denoising processes. Innovations such as Soraβs unified frameworkΒ (OpenAI, 2024) and SkyReels-V2βs infinite-length generationΒ (Chen et al., 2025a) demonstrate enhanced controllability, while efficiency gains emerge from quantizationΒ (Zhao et al., 2025), sparse attention mechanismsΒ (Xi et al., 2025; Zhang et al., 2025) and cache techniquesΒ (Wu et al., 2025; Liu et al., 2024a). The emergence of large-scale models like HunyuanVideoΒ (Kong et al., 2024) has further improved video generation quality with its causal 3D VAE architecture and full-attention mechanisms, achieving film-grade quality and physics-aware generation. These developments collectively demonstrate enhanced controllability in temporal coherence and resolution, though challenges persist in computational efficiency.
Prior work in diffusion model quantization has approached these challenges through two distinct paradigms. Quantization-Aware Training (QAT) requires full model fine-tuning but achieves superior low-bit performance. Ter-DiTΒ (Lu et al., 2024) introduces RMSNorm-enhanced adaLN modules for stable ternary training of Diffusion Transformers. In contrast, Post-Training Quantization (PTQ) methodsΒ (Li* et al., 2025; Chen et al., 2025b; Zhao et al., 2025) offer deployment-friendly solutions without retraining. SVDQuantΒ (Li* et al., 2025) employs low-rank SVD decomposition to absorb outliers into 16-bit branches for 4-bit weight/activation quantization on FLUX.1 modelsΒ (Labs, 2024). ViDiT-QΒ (Zhao et al., 2025) introduces a set of techniques for DiT-based generative models and achieves negligible performance drop under W8A8 quantization. These approaches demonstrate the trade-off between QATβs higher accuracy and PTQβs faster deployment in DiT compression.
Although PTQ offers a plug-and-play alternative, existing PTQ methods still face two critical limitations. First, the calibration-based pre-scaling adopted by most video-specific quantization techniquesΒ (Zhao et al., 2025; Wu et al., 2024; Chen et al., 2025b) not only demands extensive calibration time but also struggles to adapt to timestep-dependent scale variations in DiTs. Second, aggressive W4A4 quantization results in significant performance degradation (Fig.Β 1), with VBench metrics dropping by 27.5% to 61.3%. Our analysis reveals three key insights to overcome these limitations: (i) Weights exhibit Gaussian-like distributions (Fig.Β 3), making fixed quantization ranges suboptimal for preserving critical parameters. (ii) Activation scales vary significantly across denoising timesteps, necessitating dynamic rather than static quantization strategies. (iii) Latent feature variations exist across diffenrent denoising timesteps, enabling the possibility of adaptive bit-width allocation during online inference.
@@ 2.1 Diffusion Transformers (DiT)
Diffusion Transformers (DiTs)Β (Peebles and Xie, 2023) represent a paradigm shift in generative modeling, replacing the traditional U-NetΒ (Saharia et al., 2022; Blattmann et al., 2023; Ramesh et al., 2022) backbone with transformer architectures. This breakthrough was attributed to the self-attention mechanismβs ability to model long-range dependenciesΒ (Vaswani et al., 2023), proving especially beneficial for high-resolution synthesis.
The success of DiTs in image generation quickly extended to video synthesis. LatteΒ (Ma et al., 2025) pioneered this transition by adapting the DiT framework for temporal modeling, achieving unprecedented coherence in text-to-video generation. This was followed by SoraΒ (OpenAI, 2024), which scaled DiTs to massive parameter counts and dataset sizes, demonstrating remarkable capabilities in long video generation with complex dynamics. Open-source implementations like HunyuanVideoΒ (Kong et al., 2024), Open-SoraΒ (Zheng et al., 2024; Peng et al., 2025) and CogvideoXΒ (Yang et al., 2024) further democratized these advancements, making them widely accessible.
Despite their impressive results, DiTs inherit fundamental challenges from both transformer architectures and diffusion processes. Attentionβs quadratic complexity is particularly problematic for video generation, where sequence length scales with spatial resolution and frame count. Moreover, the iterative nature of diffusion models requires multiple forward passes through the entire architecture (typically 50-100 steps), with each step processing increasingly refined features. These limitations have driven efforts to boost DiTsβ efficiency, such as cachingΒ (Wu et al., 2025; Liu et al., 2024a), and quantizationΒ (Li* et al., 2025; Chen et al., 2025b; Zhao et al., 2025).
@@ 2.2 Model Quantization
Model quantizationΒ (Xiao et al., 2023; Li* et al., 2025; Liu et al., 2024b; Jacob et al., 2018a; Li et al., 2023) has emerged as a fundamental technique for compressing deep neural networks by converting full-precision parameters into low-bit representations, significantly reducing memory footprint and computational costs. Post-Training Quantization (PTQ)Β (Li et al., 2025; Frantar et al., 2023; Yan et al., 2025) has proven particularly effective as it compresses pre-trained models without requiring extensive retraining.
Significant progress has been made to address quantization challenges in transformersΒ (Vaswani et al., 2017). SmoothQuantΒ (Xiao et al., 2023) introduces channel-wise scaling to balance weight and activation quantization difficulty, while QuarotΒ (Ashkboos et al., 2024) employs orthogonal matrix rotations to distribute values evenly across channels. For diffusion models, quantization presents unique challenges due to their time-dependent nature. Q-DiffusionΒ (Li et al., 2023) and PTQ4DMΒ (Shang et al., 2023) address this by collecting timestep-wise activation statistics to determine optimal quantization parameters. Q-DiTΒ (Chen et al., 2025b) tackles channel-wise imbalance via customized quantization parameters. Subsequent work such as SVDQuantΒ (Li* et al., 2025) introduces low-rank branches to handle activation outliers, enabling 4-bit quantization without performance degradation. For video generation, ViDiT-QΒ (Zhao et al., 2025) achieves lossless W8A8 quantization.
Despite these advances, text-to-video generation with quantized DiTs still suffers from two persistent limitations. First, offline calibration for activation pre-scaling in existing video quantization methodsΒ (Zhao et al., 2025; Wu et al., 2024; Chen et al., 2025b) imposes heavy computational burdens and struggles to adapt to large activation vatiants. Second, pushing quantization precision below 8 bits for activations (W4A4) triggers significant quality degradation. Our framework overcomes these obstacles through three key innovations presented in Sec.Β 3.
@@ 3 Method
Overview.β As shown in Fig.Β 2, Bounded-init Grid Refinement (BGR) iteratively refines quantization grid upon progressive tightening bounds, which preserves Gaussian-distributed DiT weights with minimal error (Sec.Β 3.1). Auto-scaling Rotated Quantization (ARQ) eliminates calibration overhead by jointly optimizing rotation and online scaling for activation outliers, which also offers greater flexibility for timestep-dependent scale variants (Sec.Β 3.2). Additionally, -Guided Bit Switching further allocates bit-widths adaptively across timesteps by tracking feature evolution (Sec.Β 3.3). The synergy of these techniques achieves lower bit-widths (e.g., W4A4 and W4A6) with negligible quality degradation, overcoming limitations of prior quantization approaches.
@@ 3.1 Bounded-init Grid Refinement (BGR)
For weight matrix , vanilla MinMax (per-channel) quantizationΒ (Jacob et al., 2018b) directly adopts the extreme values for range calculation through Quant/DeQuant processes:
@@
Notably, the heuristic solution from MinMax algorithm provides an initialization for this optimization problem. As post-training quantization aims to avoid computationally expensive process (e.g., gradient descent), we seek a closed-form solution for the quantization objective. Assuming we have obtained the initial state of , and , our goal is to refine one or all of them to reduce quantization error. Refining them simultaneously is highly challenging, instead, we can refine one parameter while fixing the other two. For instance, we first fix and , then derive a better solution for by solving the least squares problem for Eq.Β 3:
@@
notably, this refinement of aligns with its definition in Eq.Β 1.
This sequential refinement procedure can be further extended to multiple rounds and combined with various initialization algorithms. Beyond MinMax, search bound methodsΒ (Liu et al., 2024b; Shao et al., 2023) also serve as valid initial strategies, which exclude outliers by progressively clipping full-precision weights:
where and denote the shrink step sizes for lower and upper bounds, respectively. Since search bound algorithm provides a better initial state for quantization step-size and zero-point, our grid refinement algorithm is built on this initialization (see Alg.Β 1).
Our experiments in Fig.Β 3 quantitatively show a substantial 86% reduction in quantization error, effectively preserving the critical parameters within high-density regions. These gains incur no additional inference-time computational overhead, making BGR more accurate than previous approaches.
@@ 3.2 Auto-scaling Rotated Quantization (ARQ)
Pre-scaling methodsΒ (Xiao et al., 2023; Zhao et al., 2025; Wu et al., 2024) introduce a channel-wise mask to alleviate quantization difficulty by transferring it from activations to weights:
where is a hyperparameter that controls the degree of difficulty mitigation. Moreover, these methods typically require a calibration dataset to compute scaling factors offline. When applied to DiT models (which involve 50 denoising stepsΒ (Xiao et al., 2023; Zhao et al., 2025)), the calibration set often fails to capture the full dynamic range of activation distributions across all timesteps, introducing substantial quantization errors and potential bias.
Rotation-based methods (Ashkboos et al., 2024; Liu et al., 2025; Zhao et al., 2025) employ an orthogonal rotation matrix satisfying and . While multiplying by can effectively suppress large outliers, the rotation operation may inadvertently amplify certain activation values. This risks introducing new quantization errors in the transformed space, as shown in Fig.Β 4.
To address these challenges, we propose Auto-scaling Rotated Quantization (ARQ) for activation quantization. Our approach combines the strengths of both rotation-based and scaling-based methods while overcoming their individual limitations. We first incorporate Hadamard matrix multiplicationΒ (Tseng et al., 2024) on both sides of activations and weights to preserve computational invariance: . Hadamard rotation is computed by fast Hadamard transform, which introduces marginal overhead to the inference latency. Subsequently, per-channel scaling factors are computed online and applied directly to activations (rather than being transferred to weights):
@@
We analyze how ARQ addresses the massive outliers and channel-wise outliers inherent in DiT activations. For massive outliers, rotation redistributes these extreme values across multiple channels. For channel-wise outliers, the post-rotation online scaling reinforces channel-wise consistency and mitigates the limitation of rotation shown in Fig.Β 4. Furthermore, our online scaling strategy eliminates reliance on offline calibration datasets. For DiT models, where activation distributions vary drastically across denoising steps, this runtime adaptation ensures optimal quantization parameters at every timestep, sustaining generation quality throughout the entire diffusion process.
@@
The denoising process in video DiTs exhibits non-uniform feature evolution across timesteps. Uniform quantization wastes computational resources on these redundant timesteps while potentially compromising quality during critical transformation phases. Existing adaptive quantization methods either depend on expensive calibration analysisΒ (Zhao et al., 2025) or adopt static timestep segmentationΒ (Wu et al., 2024). Neither of them can capture dynamic feature changes with respect to different input prompts in video generation.
@@
To further enhance video generation quality, we propose -Guided Bit Switching, tailored to input characteristics. Prior workΒ (Kahatapitiya et al., 2024; Liu et al., 2024a; Wu et al., 2025; So et al., 2024; Ma et al., 2024) found that denoising process of DiTs contains redundant timesteps, where latent feature changes are marginal. For these redundant timesteps, we can apply lower bit-width quantization across the entire model. In contrast, we use higher precision for critical timesteps with significant feature transformations. Specifically, our mixed-precision mechanism operates as follows:
@@
where denotes the most recent timestep at which cumulative error tracking is reset, and represents the modelβs output features at timestep . Our algorithm works by continuously monitoring normalized L1 distances between successive outputs. When cumulative feature changes remain below the threshold , we apply -bit quantization, indicating marginal feature evolution can tolerate lower precision. Once the accumulated error exceeds , we switch to -bit quantization to preserve critical details, while simultaneously resetting the cumulative error counter to zero. This adaptive approach optimizes bit allocation while incurring negligible additional inference overhead, with the threshold acting as a hyperparameter to balance performance and efficiency.
| Method | Bit-width (W/A) | Aesthetic Quality | Imaging Quality | Overall Consist. | Scene Consist. | BG. Consist. | Subject. Consist. | Dynamic Degree | Motion Smooth. |
| HunyuanVideoΒ (Kong et al., 2024) | 16/16 | 62.53 | 64.78 | 25.86 | 42.81 | 97.01 | 96.05 | 51.39 | 99.30 |
| MinMaxΒ (Jacob et al., 2018b) | 4/8 | 59.44 | 60.62 | 25.78 | 36.41 | 97.61 | 95.83 | 52.78 | 98.89 |
| SmoothQuantΒ (Xiao et al., 2023) | 4/8 | 60.50 | 64.47 | 25.56 | 28.85 | 97.72 | 96.29 | 51.39 | 99.05 |
| QuarotΒ (Ashkboos et al., 2024) | 4/8 | 58.80 | 56.86 | 25.33 | 34.30 | 98.10 | 95.72 | 55.56 | 99.03 |
| ViDiT-QΒ (Zhao et al., 2025) | 4/8 | 57.01 | 59.74 | 24.77 | 27.11 | 97.37 | 95.16 | 48.61 | 99.06 |
| -Quant | 4/6 | 62.27 | 64.22 | 25.83 | 33.07 | 97.89 | 96.57 | 58.33 | 99.05 |
| MinMaxΒ (Jacob et al., 2018b) | 4/4 | 24.20 | 24.78 | 4.27 | 0.00 | 98.05 | 96.27 | 0.00 | 99.03 |
| SmoothQuantΒ (Xiao et al., 2023) | 4/4 | 48.41 | 59.46 | 21.09 | 7.84 | 96.72 | 94.97 | 1.39 | 98.79 |
| QuarotΒ (Ashkboos et al., 2024) | 4/4 | 44.85 | 54.30 | 17.33 | 0.94 | 97.69 | 92.64 | 87.5 | 92.22 |
| ViDiT-QΒ (Zhao et al., 2025) | 4/4 | 45.36 | 40.10 | 19.66 | 7.85 | 97.19 | 97.29 | 0.00 | 99.43 |